OpenStudy (anonymous):

Find the 34th term of the arithmetic sequence: an = 5 + (n – 1)(-10).

OpenStudy (kropot72):

\[a _{n}=5+(n-1)(-10)=5-10n+10=15-10n\] \[a _{34}=15-(10\times 34)\] Can you work out the answer now?

OpenStudy (anonymous):

Substitute \(n=34\) to obtain \(a_{34}\).